OpenStudy (anonymous):

Prove the Well ordering property of Z by induction and prove the minimal element is unique It seems so obvious lol but I don't really know how to go about it

OpenStudy (amistre64):

n < n+1 basis step: 1 < 2, is true assume it is true for some k in Z k < k+1 ; add 1 to each side ? k+1 < (k+1)+1

OpenStudy (anonymous):

I think I have to use sets though.. The books defines it as: If \(A\) is any nonempty subset of \(\mathbb{Z}^+\), there is some element \(m \in A\) such that \(m < a\), for all \(a \in A\) (\(m\) is called a minimal element of \(A\)). So my basis step was \(1 \in A\), then obviously a has a least element since \(1 \leq n\) for every \(n \in \mathbb{Z}^+\) Then assume \(m \in A\), either m is a least element of A, or there is an integer n so that \(1 \leq n < m\) And similarly the same would go for n, so you always have a least element but I'm not sure if this is a good way to prove it

OpenStudy (amistre64):

ive seen proofs that show that if p <= q, and q <= p, then p=q which is the standard practice of showing that two values p and q are the same. but i still get a headache trying to read thru them :/

OpenStudy (amistre64):

i never really know how much of the wheel has to be reinvented during "proofs" either

OpenStudy (anonymous):

But it seems so weird to prove that it's unique since Z contains no duplicate elements so any element of a subset should be unique right?

OpenStudy (anonymous):

wait scratch that brainfart

OpenStudy (anonymous):

For proving it's unique: Let \(m, n \in A\) Assume \(n \in A\) is a least element, then \(n \leq a, \quad \forall a\in A \) So \(n \leq m\) Similary, \(m \leq n\) so \(m = n\) and it's unique

OpenStudy (anonymous):

That would work right :D

OpenStudy (amistre64):

looks familiar yes :)

OpenStudy (anonymous):

Is the induction proof above (that it has a least element) good or not? I think it is but I'm not that great with proofs :s